OpenStudy (anonymous):

Determine the rational roots of 2x 5 - 11x 4 + 14x 3 - 2x 2 + 12x + 9 = 0.

OpenStudy (anonymous):

I really need help with this question. :)

OpenStudy (mertsj):

Have you made a list of the possible rational roots?

OpenStudy (anonymous):

Not yet.

OpenStudy (mertsj):

Do that and then test them using synthetic division until you find one whose remainder is 0

OpenStudy (anonymous):

Ok. I've done that step. Is the possible answer, 1/2, 1, 3/2, 3, 9/2 and 9?

OpenStudy (mertsj):

The possible rational roots are: \[\pm1, \pm3, \pm9, \pm\frac{1}{2}, \pm\frac{3}{2}, \pm\frac{9}{2}\]

OpenStudy (anonymous):

That's what I got. :) Thanks so much!

OpenStudy (mertsj):

That is NOT the answer. You do know that, right?

OpenStudy (mertsj):

The problem says to find the rational roots. Step 1 is to make a list of the possibilities. That is what we have done so far.

OpenStudy (anonymous):

Ok. What do we do next?

OpenStudy (mertsj):

Do that and then test them using synthetic division until you find one whose remainder is 0

OpenStudy (mertsj):

Do you know how to do that?

OpenStudy (anonymous):

No, not really. Can you help me with that?

OpenStudy (mertsj):